\begin{thebibliography}{10}

\bibitem{abbass2001mbo}
Hussein~A Abbass.
\newblock Mbo: Marriage in honey bees optimization-a haplometrosis polygynous
  swarming approach.
\newblock In {\em Evolutionary Computation, 2001. Proceedings of the 2001
  Congress on}, volume~1, pages 207--214. IEEE, 2001.

\bibitem{atashpaz2007imperialist}
Esmaeil Atashpaz-Gargari and Caro Lucas.
\newblock Imperialist competitive algorithm: an algorithm for optimization
  inspired by imperialistic competition.
\newblock In {\em Evolutionary Computation, 2007. CEC 2007. IEEE Congress on},
  pages 4661--4667. IEEE, 2007.

\bibitem{bastos2009fish}
Carmelo~JA Bastos~Filho, Fernando~B de~Lima~Neto, Anthony~JCC Lins,
  Ant{\^o}nio~IS Nascimento, and Mar{\'\i}lia~P Lima.
\newblock Fish school search.
\newblock In {\em Nature-Inspired Algorithms for Optimisation}, pages 261--277.
  Springer, 2009.

\bibitem{bishop1989stochastic}
JM~Bishop.
\newblock Stochastic searching networks.
\newblock In {\em Artificial Neural Networks, 1989., First IEE International
  Conference on (Conf. Publ. No. 313)}, pages 329--331. IET, 1989.

\bibitem{chen2010hierarchical}
Hanning Chen, Yunlong Zhu, Kunyuan Hu, and Xiaoxian He.
\newblock Hierarchical swarm model: a new approach to optimization.
\newblock {\em Discrete Dynamics in Nature and Society}, 2010, 2010.

\bibitem{chen2007novel}
Tai-Chen Chen, Pei-Wei Tsai, Shu-Chuan Chu, and Jeng-Shyang Pan.
\newblock A novel optimization approach: bacterial-ga foraging.
\newblock In {\em Innovative Computing, Information and Control, 2007.
  ICICIC'07. Second International Conference on}, pages 391--391. IEEE, 2007.

\bibitem{Chu2006854}
S.-A. Chu, P.-W. Tsai, and J.-S. Pan.
\newblock Cat swarm optimization.
\newblock {\em Lecture Notes in Computer Science (including subseries Lecture
  Notes in Artificial Intelligence and Lecture Notes in Bioinformatics)}, 4099
  LNAI:854--858, 2006.
\newblock cited By (since 1996) 8.

\bibitem{chu2008fast}
Ying Chu, Hua Mi, Huilian Liao, Zhen Ji, and QH~Wu.
\newblock A fast bacterial swarming algorithm for high-dimensional function
  optimization.
\newblock In {\em Evolutionary Computation, 2008. CEC 2008.(IEEE World Congress
  on Computational Intelligence). IEEE Congress on}, pages 3135--3140. IEEE,
  2008.

\bibitem{Civicioglu201358}
P.~Civicioglu.
\newblock Artificial cooperative search algorithm for numerical optimization
  problems.
\newblock {\em Information Sciences}, 229:58--76, 2013.

\bibitem{civicioglu2012transforming}
Pinar Civicioglu.
\newblock Transforming geocentric cartesian coordinates to geodetic coordinates
  by using differential search algorithm.
\newblock {\em Computers \& Geosciences}, 46:229--247, 2012.

\bibitem{civicioglu2013backtracking}
Pinar Civicioglu.
\newblock Backtracking search optimization algorithm for numerical optimization
  problems.
\newblock {\em Applied Mathematics and Computation}, 219(15):8121--8144, 2013.

\bibitem{comellas2011bumblebees}
Francesc de~Paula Comellas~Padr{\'o}, Jes{\'u}s Mart{\'\i}nez~Navarro, et~al.
\newblock Bumblebees: a multiagent combinatorial optimization algorithm
  inspired by social insect behaviour.
\newblock 2011.

\bibitem{cuevas2012circle}
Erik Cuevas, Diego Oliva, Daniel Zaldivar, Marco P{\'e}rez-Cisneros, and
  Humberto Sossa.
\newblock Circle detection using electro-magnetism optimization.
\newblock {\em Information Sciences}, 182(1):40--55, 2012.

\bibitem{de2008novel}
FB~de~Lima~Neto, AJCC Lins, Ant{\^o}nio~IS Nascimento, Mar{\'\i}lia~P Lima,
  et~al.
\newblock A novel search algorithm based on fish school behavior.
\newblock In {\em Systems, Man and Cybernetics, 2008. SMC 2008. IEEE
  International Conference on}, pages 2646--2651. IEEE, 2008.

\bibitem{dorigo1992optimization}
Marco Dorigo.
\newblock Optimization, learning and natural algorithms.
\newblock {\em Ph. D. Thesis, Politecnico di Milano, Italy}, 1992.

\bibitem{drias2005cooperative}
Habiba Drias, Souhila Sadeg, and Safa Yahi.
\newblock Cooperative bees swarm for solving the maximum weighted
  satisfiability problem.
\newblock In {\em Computational Intelligence and Bioinspired Systems}, pages
  318--325. Springer, 2005.

\bibitem{eskandar2012water}
Hadi Eskandar, Ali Sadollah, Ardeshir Bahreininejad, and Mohd Hamdi.
\newblock Water cycle algorithm--a novel metaheuristic optimization method for
  solving constrained engineering optimization problems.
\newblock {\em Computers \& Structures}, 2012.

\bibitem{Eusuff2003210}
M.M. Eusuff and K.E. Lansey.
\newblock Optimization of water distribution network design using the shuffled
  frog leaping algorithm.
\newblock {\em Journal of Water Resources Planning and Management},
  129(3):210--225, 2003.
\newblock cited By (since 1996) 297.

\bibitem{ferreira2001gene}
Candida Ferreira.
\newblock Gene expression programming: a new adaptive algorithm for solving
  problems.
\newblock {\em arXiv preprint cs/0102027}, 2001.

\bibitem{fister2013comprehensive}
Iztok Fister, Iztok Fister~Jr., Xin-She Yang, and Janez Brest.
\newblock A comprehensive review of firefly algorithms.
\newblock {\em Swarm and Evolutionary Computation}, 2013.

\bibitem{formato2007central}
Richard~A Formato.
\newblock Central force optimization: A new metaheuristic with applications in
  applied electromagnetics.
\newblock {\em Progress In Electromagnetics Research}, 77:425--491, 2007.

\bibitem{gandomi2012krill}
Amir~Hossein Gandomi and Amir~Hossein Alavi.
\newblock Krill herd: a new bio-inspired optimization algorithm.
\newblock {\em Communications in Nonlinear Science and Numerical Simulation},
  2012.

\bibitem{geem2001new}
Zong~Woo Geem, Joong~Hoon Kim, and GV~Loganathan.
\newblock A new heuristic optimization algorithm: harmony search.
\newblock {\em Simulation}, 76(2):60--68, 2001.

\bibitem{hatamlou2012black}
Abdolreza Hatamlou.
\newblock Black hole: A new heuristic optimization approach for data
  clustering.
\newblock {\em Information Sciences}, 2012.

\bibitem{havens2008roach}
Timothy~C Havens, Christopher~J Spain, Nathan~G Salmon, and James~M Keller.
\newblock Roach infestation optimization.
\newblock In {\em Swarm Intelligence Symposium, 2008. SIS 2008. IEEE}, pages
  1--7. IEEE, 2008.

\bibitem{he2009group}
S~He, QH~Wu, and JR~Saunders.
\newblock Group search optimizer: an optimization algorithm inspired by animal
  searching behavior.
\newblock {\em Evolutionary Computation, IEEE Transactions on}, 13(5):973--990,
  2009.

\bibitem{hedayatzadeh2010termite}
Ramin Hedayatzadeh, Foad Akhavan~Salmassi, Manijeh Keshtgari, Reza Akbari, and
  Koorush Ziarati.
\newblock Termite colony optimization: A novel approach for optimizing
  continuous problems.
\newblock In {\em Electrical Engineering (ICEE), 2010 18th Iranian Conference
  on}, pages 553--558. IEEE, 2010.

\bibitem{hernandez2012distributed}
Hugo Hern{\'a}ndez and Christian Blum.
\newblock Distributed graph coloring: an approach based on the calling behavior
  of japanese tree frogs.
\newblock {\em Swarm Intelligence}, 6(2):117--150, 2012.

\bibitem{iordache2010consultant}
Serban Iordache.
\newblock Consultant-guided search: a new metaheuristic for combinatorial
  optimization problems.
\newblock In {\em Proceedings of the 12th annual conference on Genetic and
  evolutionary computation}, pages 225--232. ACM, 2010.

\bibitem{jung2003queen}
Sung~Hoon Jung.
\newblock Queen-bee evolution for genetic algorithms.
\newblock {\em Electronics letters}, 39(6):575--576, 2003.

\bibitem{karaboga2007powerful}
Dervis Karaboga and Bahriye Basturk.
\newblock A powerful and efficient algorithm for numerical function
  optimization: artificial bee colony (abc) algorithm.
\newblock {\em Journal of global optimization}, 39(3):459--471, 2007.

\bibitem{kashan2009league}
Ali~Husseinzadeh Kashan.
\newblock League championship algorithm: a new algorithm for numerical function
  optimization.
\newblock In {\em Soft Computing and Pattern Recognition, 2009. SOCPAR'09.
  International Conference of}, pages 43--48. IEEE, 2009.

\bibitem{kaveh2013new}
A~Kaveh and N~Farhoudi.
\newblock A new optimization method: Dolphin echolocation.
\newblock {\em Advances in Engineering Software}, 59:53--70, 2013.

\bibitem{kaveh2010novel}
A~Kaveh and S~Talatahari.
\newblock A novel heuristic optimization method: charged system search.
\newblock {\em Acta Mechanica}, 213(3-4):267--289, 2010.

\bibitem{kennedy1995particle}
James Kennedy and Russell Eberhart.
\newblock Particle swarm optimization.
\newblock In {\em Neural Networks, 1995. Proceedings., IEEE International
  Conference on}, volume~4, pages 1942--1948. IEEE, 1995.

\bibitem{kirkpatrick1983optimization}
Scott Kirkpatrick, D.~Gelatt Jr., and Mario~P Vecchi.
\newblock Optimization by simulated annealing.
\newblock {\em Science}, 220(4598):671--680, 1983.

\bibitem{krishnanand2005detection}
KN~Krishnanand and D~Ghose.
\newblock Detection of multiple source locations using a glowworm metaphor with
  applications to collective robotics.
\newblock In {\em Swarm Intelligence Symposium, 2005. SIS 2005. Proceedings
  2005 IEEE}, pages 84--91. IEEE, 2005.

\bibitem{krishnanand2009glowworm}
KN~Krishnanand and D~Ghose.
\newblock Glowworm swarm optimisation: a new method for optimising multi-modal
  functions.
\newblock {\em International Journal of Computational Intelligence Studies},
  1(1):93--119, 2009.

\bibitem{Li200232}
X.-L. Li, Z.-J. Shao, and J.-X. Qian.
\newblock Optimizing method based on autonomous animats: Fish-swarm algorithm.
\newblock {\em Xitong Gongcheng Lilun yu Shijian/System Engineering Theory and
  Practice}, 22(11):32, 2002.

\bibitem{lucic2001bee}
P~Lucic and D~Teodorovic.
\newblock Bee system: modeling combinatorial optimization transportation
  engineering problems by swarm intelligence.
\newblock In {\em Preprints of the TRISTAN IV triennial symposium on
  transportation analysis}, pages 441--445, 2001.

\bibitem{maia2012bee}
Renato~Dourado Maia, Leandro~Nunes de~Castro, and Walmir~Matos Caminhas.
\newblock Bee colonies as model for multimodal continuous optimization: The
  optbees algorithm.
\newblock In {\em Evolutionary Computation (CEC), 2012 IEEE Congress on}, pages
  1--8. IEEE, 2012.

\bibitem{mehrabian2006novel}
A~Reza Mehrabian and C~Lucas.
\newblock A novel numerical optimization algorithm inspired from weed
  colonization.
\newblock {\em Ecological Informatics}, 1(4):355--366, 2006.

\bibitem{mozaffari2012great}
Ahmad Mozaffari, Alireza Fathi, and Saeed Behzadipour.
\newblock The great salmon run: a novel bio--inspired algorithm for artificial
  system design and optimisation.
\newblock {\em International Journal of Bio-Inspired Computation},
  4(5):286--301, 2012.

\bibitem{mucherino2007monkey}
Antonio Mucherino and Onur Seref.
\newblock Monkey search: a novel metaheuristic search for global optimization.
\newblock In {\em Data Mining, Systems Analysis and Optimization in
  Biomedicine}, volume 953, pages 162--173, 2007.

\bibitem{parpinelli2011eco}
RS~Parpinelli and HS~Lopes.
\newblock An eco-inspired evolutionary algorithm applied to numerical
  optimization.
\newblock In {\em Nature and Biologically Inspired Computing (NaBIC), 2011
  Third World Congress on}, pages 466--471. IEEE, 2011.

\bibitem{passino2002biomimicry}
Kevin~M Passino.
\newblock Biomimicry of bacterial foraging for distributed optimization and
  control.
\newblock {\em Control Systems, IEEE}, 22(3):52--67, 2002.

\bibitem{pham2006bees}
DT~Pham, A~Ghanbarzadeh, E~Koc, S~Otri, S~Rahim, and M~Zaidi.
\newblock The bees algorithm-a novel tool for complex optimisation problems.
\newblock In {\em Proceedings of the 2nd Virtual International Conference on
  Intelligent Production Machines and Systems (IPROMS 2006)}, pages 454--459,
  2006.

\bibitem{premaratne2009new}
Upeka Premaratne, Jagath Samarabandu, and Tarlochan Sidhu.
\newblock A new biologically inspired optimization algorithm.
\newblock In {\em Industrial and Information Systems (ICIIS), 2009
  International Conference on}, pages 279--284. IEEE, 2009.

\bibitem{rabanal2007using}
Pablo Rabanal, Ismael Rodr{\'\i}guez, and Fernando Rubio.
\newblock Using river formation dynamics to design heuristic algorithms.
\newblock In {\em Unconventional Computation}, pages 163--177. Springer, 2007.

\bibitem{rashedi2009gsa}
Esmat Rashedi, Hossein Nezamabadi-Pour, and Saeid Saryazdi.
\newblock Gsa: a gravitational search algorithm.
\newblock {\em Information sciences}, 179(13):2232--2248, 2009.

\bibitem{ryan1998grammatical}
Conor Ryan, JJ~Collins, and Michael~O Neill.
\newblock Grammatical evolution: Evolving programs for an arbitrary language.
\newblock In {\em Genetic Programming}, pages 83--96. Springer, 1998.

\bibitem{shah_hosseini2007problem}
Hamed Shah-Hosseini.
\newblock Problem solving by intelligent water drops.
\newblock In {\em Evolutionary Computation, 2007. CEC 2007. IEEE Congress on},
  pages 3226--3231. IEEE, 2007.

\bibitem{shah2011principal}
Hamed Shah-Hosseini.
\newblock Principal components analysis by the galaxy-based search algorithm: a
  novel metaheuristic for continuous optimisation.
\newblock {\em International Journal of Computational Science and Engineering},
  6(1):132--140, 2011.

\bibitem{shayeghi2012anarchic}
H~Shayeghi and J~Dadashpour.
\newblock Anarchic society optimization based pid control of an automatic
  voltage regulator (avr) system.
\newblock {\em Electrical and Electronic Engineering}, 2(4):199--207, 2012.

\bibitem{shi2011optimization}
Yuhui Shi.
\newblock An optimization algorithm based on brainstorming process.
\newblock {\em International Journal of Swarm Intelligence Research (IJSIR)},
  2(4):35--62, 2011.

\bibitem{simon2008biogeography}
Dan Simon.
\newblock Biogeography-based optimization.
\newblock {\em Evolutionary Computation, IEEE Transactions on}, 12(6):702--713,
  2008.

\bibitem{storn1997differential}
Rainer Storn and Kenneth Price.
\newblock Differential evolution--a simple and efficient heuristic for global
  optimization over continuous spaces.
\newblock {\em Journal of global optimization}, 11(4):341--359, 1997.

\bibitem{su2007good}
Shoubao Su, Jiwen Wang, Wangkang Fan, and Xibing Yin.
\newblock Good lattice swarm algorithm for constrained engineering design
  optimization.
\newblock In {\em Wireless Communications, Networking and Mobile Computing,
  2007. WiCom 2007. International Conference on}, pages 6421--6424. IEEE, 2007.

\bibitem{sur2013egyptian}
Chiranjib Sur, Sanjeev Sharma, and Anupam Shukla.
\newblock Egyptian vulture optimization algorithm--a new nature inspired
  meta-heuristics for knapsack problem.
\newblock In {\em The 9th International Conference on Computing and
  InformationTechnology (IC2IT2013)}, pages 227--237. Springer, 2013.

\bibitem{tamura2011spiral}
Kenichi Tamura and Keiichiro Yasuda.
\newblock Spiral dynamics inspired optimization.
\newblock {\em Journal of Advanced Computational Intelligence and Intelligent
  Informatics}, 15(8):1116--1122, 2011.

\bibitem{6360147}
Rui Tang, S.~Fong, Xin-She Yang, and S.~Deb.
\newblock Wolf search algorithm with ephemeral memory.
\newblock In {\em Digital Information Management (ICDIM), 2012 Seventh
  International Conference on}, pages 165--172, 2012.

\bibitem{teodorovic2005bee}
Du{\v{s}}an Teodorovi{\'c} and Mauro Dell'Orco.
\newblock Bee colony optimization--a cooperative learning approach to complex
  transportation problems.
\newblock In {\em Advanced OR and AI Methods in Transportation: Proceedings of
  16th Mini--EURO Conference and 10th Meeting of EWGT (13-16 September
  2005).--Poznan: Publishing House of the Polish Operational and System
  Research}, pages 51--60, 2005.

\bibitem{ting2012weightless}
TO~Ting, Ka~Lok Man, Sheng-Uei Guan, Mohamed Nayel, and Kaiyu Wan.
\newblock Weightless swarm algorithm (wsa) for dynamic optimization problems.
\newblock In {\em Network and Parallel Computing}, pages 508--515. Springer,
  2012.

\bibitem{vicsek1995novel}
Tam{\'a}s Vicsek, Andr{\'a}s Czir{\'o}k, Eshel Ben-Jacob, Inon Cohen, and Ofer
  Shochet.
\newblock Novel type of phase transition in a system of self-driven particles.
\newblock {\em Physical Review Letters}, 75(6):1226--1229, 1995.

\bibitem{Wedde200483}
H.F. Wedde, M.~Farooq, and Y.~Zhang.
\newblock Beehive: An efficient fault-tolerant routing algorithm inspired by
  honey bee behavior.
\newblock {\em Lecture Notes in Computer Science (including subseries Lecture
  Notes in Artificial Intelligence and Lecture Notes in Bioinformatics)}, 3172
  LNCS:83--94, 2004.

\bibitem{xu2010social}
Yuechun Xu, Zhihua Cui, and Jianchao Zeng.
\newblock Social emotional optimization algorithm for nonlinear constrained
  optimization problems.
\newblock In {\em Swarm, Evolutionary, and Memetic Computing}, pages 583--590.
  Springer, 2010.

\bibitem{yan2013novel}
GAO-WEI YAN and ZHAN-JU HAO.
\newblock A novel optimization algorithm based on atmosphere clouds model.
\newblock {\em International Journal of Computational Intelligence and
  Applications}, 2013.

\bibitem{Yang2005317}
X.-S. Yang.
\newblock Engineering optimizations via nature-inspired virtual bee algorithms.
\newblock volume 3562, pages 317--323, 2005.

\bibitem{yang2008book}
Xin-She Yang.
\newblock {\em Nature-Inspired Metaheuristic Algorithms}.
\newblock Luniver Press, UK, 2008.

\bibitem{yang2010firefly}
Xin-She Yang.
\newblock Firefly algorithm, stochastic test functions and design optimisation.
\newblock {\em International Journal of Bio-Inspired Computation}, 2(2):78--84,
  2010.

\bibitem{yang2010natureluniver}
Xin-She Yang.
\newblock {\em Nature-Inspired Metaheuristic Algorithms, 2nd Edition}.
\newblock Luniver Press, 2010.

\bibitem{yang2012flower}
Xin-She Yang.
\newblock Flower pollination algorithm for global optimization.
\newblock {\em Unconventional Computation and Natural Computation}, pages
  240--249, 2012.

\bibitem{yangswarmbook2013}
Xin-She Yang, Zhihua Cui, Renbin Xiao, Amir~Hossein Gandom, and Mehmet
  Karamanoglu.
\newblock {\em Swarm Intelligence and Bio-inspired Computation: Theory and
  Applications}.
\newblock Elsevier, 2013.

\bibitem{yang2009cuckoo}
Xin-She Yang and Suash Deb.
\newblock Cuckoo search via l{\'e}vy flights.
\newblock In {\em Nature \& Biologically Inspired Computing, 2009. NaBIC 2009.
  World Congress on}, pages 210--214. IEEE, 2009.

\bibitem{yang2010eagle}
Xin-She Yang and Suash Deb.
\newblock Eagle strategy using l\'evy walk and firefly algorithms for
  stochastic optimization.
\newblock In {\em Nature Inspired Cooperative Strategies for Optimization
  (NICSO2010)}, pages 101--111. Springer, 2010.

\bibitem{yangflower2013}
Xin-She Yang, Mehmet Karamanoglu, and Xingshi He.
\newblock Multi-objective flower algorithm for optimization.
\newblock {\em Procedia Computer Science}, 18:861--868, 2013.

\bibitem{yang2006application}
Xin-She Yang, Janet~M Lees, and Chris~T Morley.
\newblock Application of virtual ant algorithms in the optimization of cfrp
  shear strengthened precracked structures.
\newblock In {\em Computational Science--ICCS 2006}, pages 834--837. Springer,
  2006.

\bibitem{Bat:2010}
X.S. Yang.
\newblock A new metaheuristic bat-inspired algorithm.
\newblock {\em Nature Inspired Cooperative Strategies for Optimization (NICSO
  2010)}, pages 65--74, 2010.

\bibitem{zandi2012reactive}
Z~Zandi, E~Afjei, and M~Sedighizadeh.
\newblock Reactive power dispatch using big bang-big crunch optimization
  algorithm for voltage stability enhancement.
\newblock In {\em Power and Energy (PECon), 2012 IEEE International Conference
  on}, pages 239--244. IEEE, 2012.

\bibitem{zhang2009human}
Luna~Mingyi Zhang, Cheyenne Dahlmann, and Yanqing Zhang.
\newblock Human-inspired algorithms for continuous function optimization.
\newblock In {\em Intelligent Computing and Intelligent Systems, 2009. ICIS
  2009. IEEE International Conference on}, volume~1, pages 318--321. IEEE,
  2009.

\end{thebibliography}

\documentclass[journal,a4paper,twoside]{sty/IEEEtran}

\usepackage [cp1250] {inputenc}
\usepackage{sty/EVjour}

\ifCLASSINFOpdf
   \usepackage[pdftex]{graphicx}
\else 
   \usepackage[dvips]{graphicx}
\fi
\usepackage{epstopdf} 

\usepackage{textcomp} 

\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{latexsym}
\usepackage{graphicx}
\usepackage{lscape}
\usepackage{multirow}
\usepackage{rotating}
\usepackage{amsmath}
\usepackage{color}

\begin{document}

\title{A Brief Review of Nature-Inspired Algorithms for Optimization}

\authors{ Iztok Fister Jr.${}^{1}$, Xin-She Yang${}^{2}$, Iztok Fister${}^{1}$, Janez Brest${}^{1}$, Du\v{s}an Fister${}^{1}$}

\address{$^1$University of Maribor
\\ Faculty of electrical engineering and computer science, Smetanova 17, 2000 Maribor, Slovenia
\\ E-mail: iztok.fister@guest.arnes.si
\\$^2$University of Middlesex
\\ School of Science and Technology, Middlesex University, London NW4 4BT, United Kingdom
\\ E-mail: xy227@cam.ac.uk}

\abstract{Swarm intelligence and bio-inspired algorithms form a hot topic in the developments of new algorithms inspired by nature. These nature-inspired metaheuristic algorithms can be based on swarm intelligence, biological systems, physical and chemical systems. Therefore, these algorithms can be called swarm-intelligence-based, bio-inspired, physics-based and chemistry-based, depending on the sources of inspiration. Though not all of them are efficient, a few algorithms have proved to be very efficient and thus have become popular tools for solving real-world problems. Some algorithms are insufficiently studied.
The purpose of this review is to present a relatively comprehensive list of all the algorithms in the literature, so as to inspire further research.}

\keywords{swarm intelligence, bio-inspired algorithms, physics/chemistry algorithms, optimization}

\review{15 July 2013}    

\markboth{Fister, Yang Fister, Brest, Fister}{A Brief Review of Nature-Inspired Optimization Algorithms}

\maketitle

\IEEEpeerreviewmaketitle

\vspace{2cm}

\section{Introduction}

Real-world optimization problems are often very challenging to solve, and many applications have to deal with NP-hard problems. To solve such problems,
optimization tools have to be used, though there is no guarantee that the optimal solution can be obtained. In fact, for NP-problems, there are no
efficient algorithms at all. As a result, many problems have to be solved by trial and errors using various optimization techniques.  In addition, new algorithms have been developed to see if they can cope with these challenging optimization problems.

Among these new algorithms, many algorithms such as particle swarm optimization, cuckoo search and firefly algorithm,
have gained popularity due to their high efficiency. In the current literature, there are about 40 different algorithms.
It is really a challenging task to classify these algorithms systematically. Obviously, the classifications can largely
depend on the criteria, and there is no easy guideline to set out the criteria in the literature.
As criteria may vary, detailed classifications can be an impossible task for a research paper.
However, in this short paper, we only attempt to focus on one aspect of the characteristics of
these algorithms. That is, we will focus on the source of inspiration when developing algorithms.

Therefore, the rest of this paper is organized as follows: Section 2 analyzes the sources
of inspiration, while Section 3 provides a brief and yet comprehensive list
of algorithms. Finally, Section 4 concludes with some suggestions.

\section{Sources of Inspiration}

Nature has inspired many researchers in many ways and thus is a rich source of inspiration.
Nowadays, most new algorithms are nature-inspired, because they have been developed
by drawing inspiration from nature. Even with the emphasis on the source of inspiration,
we can still have different levels of
classifications, depending on how details and how many subsources we will wish to use.
For simplicity, we will use the highest level sources such as biology, physics or chemistry.

In the most generic term, the main source of inspiration is Nature. Therefore, almost all
new algorithms can be referred to as nature-inspired. By far the majority of nature-inspired
algorithms are based on some successful characteristics of biological system. Therefore, the
largest fraction of nature-inspired algorithms are biology-inspired, or bio-inspired for short.

Among bio-inspired algorithms, a special class of algorithms have been developed by drawing
inspiration from swarm intelligence. Therefore, some of the bio-inspired algorithms can be
called swarm-intelligence-based. In fact, algorithms based on swarm intelligence are among the
most popular. Good examples are ant colony optimization \cite{dorigo1992optimization},
particle swarm optimization \cite{kennedy1995particle},
cuckoo search \cite{yang2009cuckoo},  bat algorithm \cite{Bat:2010}, and firefly algorithm \cite{yang2008book,fister2013comprehensive}.

Obviously, not all algorithms were based on biological systems. Many algorithms have been
developed by using inspiration from physical and chemical systems. Some may even be based on
music \cite{geem2001new}. In the rest of paper, we will briefly divide all algorithms into
different categories, and we do not claim that this categorization is unique. This is a good attempt
to provide sufficiently detailed references.

\section{Classification of Algorithms}

Based on the above discussions, we can divide all existing algorithms into four major
categories: swarm intelligence (SI) based, bio-inspired (but not SI-based), physics/chemistry-based,
and others. We will summarize them briefly in the rest of this paper.
However, we will focus here on the relatively new algorithms. Well-established algorithms such as genetic algorithms
are so well known that there is no need to introduce them in this brief paper.

It is worth pointing out the classifications here are not unique as some algorithms can be classified into
different categories at the same time. Loosely speaking, classifications depend largely on
what the focus or emphasis and the perspective may be. For example, if the focus and perspective
are about the trajectory of the search path, algorithms can be classified as trajectory-based
and population-based. Simulated annealing is a good example of trajectory-based algorithms,
while particle swarm optimization and firefly algorithms are population-based algorithms.
If our emphasis is placed on the interaction of the multiple agents, algorithms can be
classified as attraction-based or non-attraction-based. Firefly algorithm (FA) is a good example
of attraction-based algorithms because FA uses the attraction of light and attractiveness
of fireflies, while genetic algorithms are non-attraction-based since there is no explicit
attraction used. On the other hand, if the emphasis is placed on the updating equations,
algorithms can be divided into rule-based and equation-based. For example, particle swarm optimization
and cuckoo search are equation-based algorithms because both use explicit updating equations,
while genetic algorithms do not have explicit equations for crossover and mutation.
However, in this case, the classifications are not unique. For example, firefly algorithm
uses three explicit rules and these three rules can be converted explicitly into a single
updating equation which is nonlinear \cite{yang2008book,fister2013comprehensive}. This clearly
shows that classifications depend on the actual perspective and motivations.
Therefore, the classifications here are just
one possible attempt, though the emphasis is placed on the sources of inspiration.

\subsection{Swarm intelligence based}

Swarm intelligence (SI) concerns the collective, emerging behaviour of multiple, interacting agents
who follow some simple rules. While each agent may be considered as unintelligent, the
whole system of multiple agents may show some self-organization behaviour and thus can behave like
some sort of collective intelligence. Many algorithms have been developed by drawing inspiration
from swarm-intelligence systems in nature.

All SI-based algorithms use multi-agents, inspired by the collective behaviour of social insects, like ants, termites, bees, and wasps, as well as from other animal societies like flocks of birds or fish. A list of swarm intelligence algorithms is presented in Table~\ref{Table1}.
The classical particle swarm optimization (PSO) uses the swarming behaviour of fish and birds,
while firefly algorithm (FA) uses the flashing behaviour of swarming fireflies.
Cuckoo search (CS) is based on the brooding parasitism of some cuckoo species,
while bat algorithm uses the echolocation of foraging bats. Ant colony optimization
uses the interaction of social insects (e.g., ants), while the class of bee algorithms are all
based on the foraging behaviour of honey bees.

SI-based algorithms are among the most popular and widely used. There are many reasons for such popularity, one of the reasons is that SI-based algorithms usually sharing information among multiple agents, so that self-organization, co-evolution and learning during iterations may help to provide the high efficiency of most SI-based algorithms. Another reason is that multiple agent can be parallelized easily so that large-scale optimization becomes more practical
from the implementation point of view.

\subsection{Bio-inspired, but not SI based}

Obviously, SI-based algorithms belong to a wider class of algorithms, called bio-inspired algorithms.
In fact, bio-inspired algorithms form a majority of all nature-inspired algorithms.
From the set theory point of view, SI-based algorithms are a subset of bio-inspired algorithms,
while bio-inspired algorithms are a subset of nature-inspired algorithms.
That is
\[ \textrm{SI-based} \subset \textrm{bio-inspired} \subset \textrm{nature-inspired}. \]
Conversely, not all nature-inspired algorithms are bio-inspired, and some are purely physics and chemistry based algorithms
as we will see below.

Many bio-inspired algorithms do not use directly the swarming behaviour.
Therefore, it is better to call them bio-inspired, but not SI-based.
For example, genetic algorithms are bio-inspired, but
not SI-based. However, it is not easy to classify certain algorithms such as differential evolution (DE). Strictly speaking,
DE is not bio-inspired because there is no direct link to any biological behaviour. However, as it has some similarity
to genetic algorithms and also has a key word `evolution', we tentatively put it in the category of bio-inspired algorithms.
These relevant algorithms are listed in Table~\ref{Table1}.

For example, the flower algorithm \cite{yang2012flower}, or flower pollination algorithm \cite{yangflower2013}, 
developed by Xin-She Yang in 2012 is a bio-inspired algorithm, but it is not a SI-based algorithm because
flower algorithm tries to mimic the pollination characteristics of flowering plants and the associated
flower consistency of some pollinating insects.

\subsection{Physics and Chemistry Based}

Not all metaheuristic algorithms are bio-inspired, because their sources of inspiration 
often come from physics and chemistry.
For the algorithms that are not bio-inspired, most have been developed by mimicking certain
physical and/or chemical laws, including electrical charges, gravity, river systems, etc.
As different natural systems are relevant to this category, we can even subdivide
these into many subcategories which is not necessary.
A list of these algorithms is given in Table~\ref{Table1}.

Schematically, we can represent the relationship of physics and chemistry based algorithms
as the follows:
\[
\frac{\textrm{Physics algorithms}}{\text{Chemistry algorithms}} \left\{ \begin{array}{l}
 \notin  \textrm{bio-inspired algorithms} \\ \\
 \in  \text{nature-inspired algorithms}
\end{array} \right.
\]
Though physics and chemistry are two different subjects, however, it is not useful to
subdivide this subcategory further into physics-based and chemistry. After all, many
fundamental laws are the same. So we simply group them as physics and chemistry based algorithms.

\subsection{Other algorithms}
When researchers develop new algorithms, some may look for inspiration away from nature.
Consequently, some algorithms are not bio-inspired or physics/chemistry-based,  it 
is sometimes difficult to put some algorithms in the above three categories, because these
algorithms have been developed by using various characteristics from different sources, such as
social, emotional, etc. In this case, it is better to put them in the other category, as listed
in Table~\ref{Table1}

\subsection{Some Remarks}
Though the sources of inspiration are very diverse, the algorithm designed from such
inspiration may be equally diverse. However, care should be taken, as true novelty is
a rare thing. For example, there are about 28,000 living species of fish,
this cannot mean that researchers should develop 28000 different algorithms based on fish.
Therefore, one cannot call their algorithms trout algorithm, squid algorithm, ..., shark algorithm.

In essence, researchers try to look for some efficient formulas as summarized by Yang
\cite{yangswarmbook2013} as the following generic scheme:
\[ [x_1, x_2, ..., x_n]^{t+1}=A\big\{[x_1, x_2, ..., x_n]^t; ...; \]
\[ \qquad (p_1,p_2, ..., p_k); (w_1, w_2, ..., w_m)  \big\}, \]
which attempts to generate better solutions (a population of $n$ solutions) at iteration $t+1$ from
the current iteration $t$ and its solution set $x_i, (i=1,2,...,n)$. This iterative algorithmic engine
(i.e., algorithm $A$) also uses some algorithm-dependent parameters $(p_1, ..., p_k)$
and some random variables $(w_1, ..., w_m)$. This schematic representation can include
all the algorithms listed in this paper. However, this does not mean it is easy
to analyze the behaviour of an algorithm because this formula can be highly nonlinear.
Though Markov chains theory and dynamical system theory can help to
provide some limited insight into some algorithms, the detailed mathematical framework
is still yet to be developed. 

On the other hand, it is worth pointing out that studies show that some algorithms are better
than others. It is still not quite understood why. However, if one looks at the intrinsic part
of algorithm design closely, some algorithms are badly designed, which lack certain basic capabilities
such as the mixing and diversity among the solutions. In contrast, good algorithms have
both mixing and diversity control so that the algorithm can explore the vast search
space efficiently, while converge relatively quickly when necessary.
Good algorithms such as particle swarm optimization, differential evolution, cuckoo search 
and firefly algorithms all have both global search and intensive local search capabilities,
which may be partly why they are so efficient.

\section{Conclusion}
The sources of inspiration for algorithm development are very diverse, and consequently the algorithms are equally diverse.
In this paper, we have briefly summarized all the algorithms into 4 categories. This can be a comprehensive
source of information to form a basis or starting point for further research. It is worth pointing out that
the classifications may not be unique, and this present table is just for the purpose of information only.

Based on many studies in the literature, some algorithms are more efficient and popular than others. It would be helpful
to carry out more studies, but this does not mean that
we should encourage researchers  to develop more algorithms such as grass, sky, or ocean algorithms.

Currently, there may be some confusion and distraction in the research
of metaheuristic algorithms. On the one hand, researchers have focused
on important novel ideas for solving difficult problems. On the other hand,
some researchers artificially invent new algorithms for the sake of
publications with little improvement and no novelty.
Researchers should be encouraged to carry out truly novel and important
studies that are really useful to solve hard problems. Therefore, our
aim is to inspire more research to gain better insight into efficient
algorithms and solve large-scale real-world problems.

\begin{landscape}
\begin{table}
\begin{tabular}{ |l|l|l|l|l|l|l|l|l| }
\hline
\multicolumn{3}{ |c| }{\textcolor{blue}{Swarm intelligence based algorithms}} & \multicolumn{3}{ |c| }{\textcolor{blue}{Bio-inspired (not SI-based) algorithms}}\\
\hline
 Algorithm & Author & Reference & Algorithm & Author & Reference \\ \hline
   Accelerated PSO & Yang et al. & \cite{yang2008book,yang2010natureluniver} & Atmosphere clouds model & Yan and Hao & \cite{yan2013novel} \\
   Ant colony optimization & Dorigo& \cite{dorigo1992optimization} & Biogeography-based optimization & Simon & \cite{simon2008biogeography} \\
   Artificial bee colony & Karaboga and Basturk & \cite{karaboga2007powerful} & Brain Storm Optimization & Shi & \cite{shi2011optimization}\\
   Bacterial foraging & Passino & \cite{passino2002biomimicry} & Differential evolution & Storn and Price & \cite{storn1997differential}\\
   Bacterial-GA Foraging & Chen et al. & \cite{chen2007novel} & Dolphin echolocation & Kaveh and Farhoudi & \cite{kaveh2013new}\\
   Bat algorithm & Yang & \cite{Bat:2010} & Japanese tree frogs calling & Hern{\'a}ndez and Blum & \cite{hernandez2012distributed}\\     	
   Bee colony optimization & Teodorovi{\'c} and Dell'Orco & \cite{teodorovic2005bee} & Eco-inspired evolutionary algorithm & Parpinelli and Lopes & \cite{parpinelli2011eco}\\
   Bee system & Lucic and Teodorovic & \cite{lucic2001bee} & Egyptian Vulture & Sur et al. & \cite{sur2013egyptian}\\
   BeeHive & Wedde et al. & \cite{Wedde200483} & Fish-school Search & Lima et al. & \cite{de2008novel,bastos2009fish}\\
    Wolf search & Tang et al. & \cite{6360147} & Flower pollination algorithm & Yang & \cite{yang2012flower,yangflower2013}\\
   Bees algorithms & Pham et al. & \cite{pham2006bees} & Gene expression & Ferreira & \cite{ferreira2001gene}\\
   Bees swarm optimization & Drias et al. & \cite{drias2005cooperative} & Great salmon run & Mozaffari & \cite{mozaffari2012great}\\
   Bumblebees & Comellas and Martinez & \cite{comellas2011bumblebees} & Group search optimizer & He et al. & \cite{he2009group}\\
   Cat swarm & Chu et al. & \cite{Chu2006854} & Human-Inspired Algorithm & Zhang et al. & \cite{zhang2009human} \\
   Consultant-guided search & Iordache & \cite{iordache2010consultant} & Invasive weed optimization & Mehrabian and Lucas & \cite{mehrabian2006novel}\\
   Cuckoo search & Yang and Deb & \cite{yang2009cuckoo} & Marriage in honey bees & Abbass & \cite{abbass2001mbo}\\
   Eagle strategy & Yang and Deb & \cite{yang2010eagle} & OptBees & Maia et al. & \cite{maia2012bee}\\
   Fast bacterial swarming algorithm & Chu et al. & \cite{chu2008fast} & Paddy Field Algorithm & Premaratne et al. & \cite{premaratne2009new}\\
   Firefly algorithm & Yang & \cite{yang2010firefly} & Roach infestation algorithm& Havens & \cite{havens2008roach}\\
   Fish swarm/school & Li et al. & \cite{Li200232} & Queen-bee evolution & Jung & \cite{jung2003queen}\\
   Good lattice swarm optimization & Su et al. & \cite{su2007good}  & Shuffled frog leaping algorithm & Eusuff and Lansey & \cite{Eusuff2003210}\\
   Glowworm swarm optimization & Krishnanand and Ghose & \cite{krishnanand2005detection,krishnanand2009glowworm} & Termite colony optimization & Hedayatzadeh et al. & \cite{hedayatzadeh2010termite} \\
   Hierarchical swarm model & Chen et al. & \cite{chen2010hierarchical} & \multicolumn{3}{ |c| }{\textcolor{blue}{Physics and Chemistry based algorithms}}  \\
   Krill Herd & Gandomi and Alavi & \cite{gandomi2012krill} & Big bang-big Crunch & Zandi et al. & \cite{zandi2012reactive}\\
    Monkey search & Mucherino and Seref & \cite{mucherino2007monkey} &  Black hole & Hatamlou & \cite{hatamlou2012black}\\
   Particle swarm algorithm & Kennedy and Eberhart & \cite{kennedy1995particle} & Central force optimization & Formato & \cite{formato2007central}\\
   Virtual ant algorithm & Yang & \cite{yang2006application} & Charged system search & Kaveh and Talatahari & \cite{kaveh2010novel}\\
   Virtual bees & Yang & \cite{Yang2005317} & Electro-magnetism optimization & Cuevas et al. & \cite{cuevas2012circle}\\
  Weightless Swarm Algorithm & Ting et al. & \cite{ting2012weightless} & Galaxy-based search algorithm & Shah-Hosseini & \cite{shah2011principal}\\
   \multicolumn{3}{ |c| }{\textcolor{blue}{Other algorithms}} & Gravitational search & Rashedi et al. & \cite{rashedi2009gsa}\\
    Anarchic society optimization & Shayeghi and Dadashpour & \cite{shayeghi2012anarchic} & Harmony search & Geem et al. & \cite{geem2001new}\\
   Artificial cooperative search & Civicioglu & \cite{Civicioglu201358}  & Intelligent water drop & Shah-Hosseini & \cite{shah_hosseini2007problem}\\
   Backtracking optimization search & Civicioglu & \cite{civicioglu2013backtracking} & River formation dynamics & Rabanal et al. & \cite{rabanal2007using}\\
   Differential search algorithm & Civicioglu & \cite{civicioglu2012transforming} & Self-propelled particles & Vicsek & \cite{vicsek1995novel}\\
    Grammatical evolution & Ryan et al. & \cite{ryan1998grammatical} & Simulated annealing & Kirkpatrick et al. & \cite{kirkpatrick1983optimization}\\
  Imperialist competitive algorithm & Atashpaz-Gargari and Lucas & \cite{atashpaz2007imperialist} & Stochastic difusion search & Bishop & \cite{bishop1989stochastic}\\
   League championship algorithm & Kashan & \cite{kashan2009league} & Spiral optimization & Tamura and Yasuda & \cite{tamura2011spiral}\\
   Social emotional optimization & Xu et al. & \cite{xu2010social} & Water cycle algorithm & Eskandar et al. & \cite{eskandar2012water} \\
\hline
\end{tabular}
\caption{A list of algorithms}
\label{Table1}
\end{table}
\end{landscape}
\normalsize

\small
\bibliographystyle{plain}
\bibliography{bibtex}
\normalsize
\vspace{0cm}

\small

\begin{IEEEbiographynophoto}{Iztok Fister Jr.} received his B.Sc. degree in Computer Science in 2011 from the Faculty of Electrical Engineering and Computer Science of the University of Maribor, Slovenia. Currently, he is working towards his M.Sc. degree. His research activities encompasses swarm intelligence, pervasive computing and programming languages.
\end{IEEEbiographynophoto}

\begin{IEEEbiographynophoto}{Xin-She Yang}
received his Ph.D degree in Applied Mathematics from
University of Oxford, UK. Now he is a Reader in Modelling and
Simulation at Middlesex University, UK, an Adjunct Professor at
Reykjavik University, Iceland, and a Distuiguished Guest Professor
at Xi'an Polytechnic University, China. He is also the IEEE CIS
Chair for the Task Force on Business Intelligene and Knowledge
Management, and the Editor-in-Chief of International Journal of
Mathematical Modelling and Numerical Optimisation (IJMMNO).
\end{IEEEbiographynophoto}

\begin{IEEEbiographynophoto}{Iztok Fister} graduated in computer science from the University of
Ljubljana in 1983. He received his Ph.D degree from the Faculty of
Electrical Engineering and Computer Science, University of Maribor,
in 2007. He works as an assistant in the Computer Architecture
and Languages Laboratory at the same faculty. His research interests
include program languages, operational researches and evolutionary
algorithms.
\end{IEEEbiographynophoto}

\begin{IEEEbiographynophoto}{Janez Brest} received his B.S., M.Sc, and Ph.D degrees in computer
science from the University of Maribor, Maribor, Slovenia, in 1995,
1998 and 2000, respectively. He is currently a full professor at the
Faculty of Electrical Engineering and Computer Science, University
of Maribor.
\end{IEEEbiographynophoto}

\begin{IEEEbiographynophoto}{Du\v{s}an Fister} is a student of the Faculty of Electrical Engineering and Computer Science of the University of Maribor. His research activities encompasses GPS solutions, swarm intelligence and operating systems.
\end{IEEEbiographynophoto}

\vfill

\label{finish}
\end{document}